### v2

@@ Acknowledgements.
Appendix A LMC rest frame
A.1 CM rest frame
The coordinate system used to compare the data with the theoretical model has its origin in the LMC and SMC centre of mass and the $x$-axis orientated according to the DM subhalo velocity $\bar{v}_{s}$. In order to obtain the coordinates of our data set in such rest frame, we have performed the following steps
@@ A.1 CM rest frame
to obtain the coordinate $\left(x^{\prime},\,y^{\prime},\,z^{\prime}\right)$ in the new rest frame. The velocity has to be transformed as well and, the final velocity must be boosted in $\bar{v}_{s}=v_{s}\hat{i}^{\prime}$. The angles are defined as
@@
where $\left(\left(v_{cm}\right)_{x},\,\left(v_{cm}\right)_{y},\,\left(v_{cm}\right)_{z}\right)$ is the centre of mass velocity in the solar coordinate system.
@@ A.2 Orbit frame
The coordinate system used to compare our results with the Fig. 1 presented by Buschmann et al. (2018) has its origin in the CM of the Magellanic Clouds System, its $z^{*}$-axis perpendicular to the orbital plane, and the $x^{*}$-axis orientated in the direction of the velocity of the DM subhalo. To obtain these new coordinates from the CM rest frame, we have to perform a rotation of an angle $\theta_{orb}$ according to the following matrix
@@
where $\tan\theta_{orb}=\overline{z}_{orb}/\overline{y}_{orb}$, $\overline{z}_{orb}$ $\left(\overline{y}_{orb}\right)$ is the mean value of the $z$ $\left(y\right)$ coordinate in the CM rest frame of the CM past orbit.
@@ Appendix B Estimation of radial velocities using Machine Learning
To complete the phase information for all the halo stars in our sample, we applied two machine learning techniques for assigning the radial velocities to those stars without such measurements.
@@ B.1 K-Giant radial velocity
Our first sample consists of 245086 K-Giant stars, where 10989 have measured radial velocities. The last subsample was used to train a Random Forest Regressor (RF) (Breiman, 2001). The chosen predictors for this study were the angular coordinates, proper motions, G magnitude, BP and RP colours, distances to the Sun and the galactic centre, and the Galactocentric Cartesian Coordinates. To prevent overfitting, a standard cross-validation analysis was performed. The RF hyper-parameters were tuned using GridSearchCV from the Scikit Learn library, resulting in the following values: [$n\_estimators=4900$, $random\_state=0$]
B.2 RRLyrae radial velocity
We observed a decrease in the number of measured radial velocities for RR Lyrae stars, resulting in a drop in the fraction of measured radial velocities to (5510/67276), considering galactocentric distances between 10 to 100 kpc. To tackle this issue, we aimed to model the spatial distribution of radial velocities. To achieve this, we utilised Normalizing Flow (NFs) (Durkan et al., 2019), as implemented by Crenshaw et al. (2021), to model the joint posterior probability of radial velocities and predictors using a subsample of the features described earlier. However, we limited the feature space to prevent any bias against less luminous stars. Therefore, we did not consider the magnitude and colours of stars as predictors in this case.
@@ B.3 Goodness of fit in radial velocity regression
R-Squared ($R^{2}$), or the coefficient of determination, is a statistical measure used to determine the proportion of variance in the dependent variable that can be explained by the independent variable in a regression model. The statistics were calculated for our two samples of stars resulting in values of 0.86 for K-Giant and 0.5 for RR-Lyrae stars respectively. These values are comparable to those recently reported by Naik & Widmark (2023) using Bayesian Neural Networks, and provide us with a complete sample of halo stars in phase space.
@@
Using this method, we can estimate the mean velocity field in the halo, which is necessary for the likelihood estimation of the subhalo mass of the Large Magellanic Cloud.
@@ Appendix C Validation of estimated distances for K-Giant stars
To test the photometric distance obtained for the K-Giants we have computed the photometric distance of the globular clusters NGC7006, NGC5694, NGC2419, NGC6229, and for the LMC, SMC, Carina, Draco, and Sculptor.
The observational data was extracted from the Gaia Data Release 3 (Gaia Collaboration et al., 2022, 2016), using the option single object searcher to avoid field-contamination. Additionally, the data for the LMC and the SMC was selected on a circular disk in $(l,b)$ centred in the location of each MC, of radius $2^{\circ}$. For Carina, Draco and Sculptor we selected data using an angular mask of 3 times the tidal radius reported in Drlica-Wagner et al. (2020a). We have performed the data reduction/treatment indicated in the text for the K-Giant (that is corrected due to dust extinction, discarded all sources with $E(B-V)>0.3$, corrected the magnitudes, and performed the $3\sigma$ cut upon the corrected BP and RP flux excess factor ($C^{*}$)). Finally, we selected the giant branch and performed the cross-match with the K-Giant catalogue given by GAIA (gaiadr3.astrophysical parameters). For the LMC and SMC, we have also restricted the data set in the reported Gaia’s parallax.
We calculated the photometric distance for each source and the mean value for each object. The results are shown in Fig. 5, and, as can be seen, the computed photometric distance and the values reported in the literature (Harris, 1996b; Drlica-Wagner et al., 2020b) are in perfect agreement.
Figure 5: Comparison between the photometric distance computed using Eq. (1) and the tabulated distance (Harris, 1996b; Drlica-Wagner et al., 2020b). Olive solid line: photometric distance equals tabulated distance; dashed-brown line: ±$\pm 10\%$ respect to the one-to-one line; purple-solid line: ±$\pm 20\%$ respect to the one-to-one line. The vertical lines are the statistical errors.

### v3

@@ Acknowledgements.
Appendix A Large Magellanic Cloud rest frame
A.1 Centre-of-mass rest frame
The coordinate system used to compare the data with the theoretical model has its origin in the LMC and SMC CM and the $x$ axis orientated according to the DM subhalo velocity, $\bar{v}_{s}$. In order to obtain the coordinates of our dataset in such a rest frame, we performed the following steps:
@@ A.1 Centre-of-mass rest frame
to obtain the coordinate $\left(x^{\prime},\,y^{\prime},\,z^{\prime}\right)$ in the new rest frame. The velocity had to be transformed as well and the final velocity had to be boosted using $\bar{v}_{s}=v_{s}\hat{i}^{\prime}$. The angles are defined as
@@
where $\left(\left(v_{cm}\right)_{x},\,\left(v_{cm}\right)_{y},\,\left(v_{cm}\right)_{z}\right)$ is the CM velocity in the solar coordinate system.
@@ A.2 Orbit frame
The coordinate system used to compare our results with the Fig. 1 presented by Buschmann et al. (2018) has its origin in the CM of the MCS, its $z^{*}$ axis perpendicular to the orbital plane, and the $x^{*}$ axis orientated in the direction of the velocity of the DM subhalo. To obtain these new coordinates from the CM rest frame, we had to perform a rotation of an angle, $\theta_{orb}$, according to the following matrix:
@@
where $\tan\theta_{orb}=\overline{z}_{orb}/\overline{y}_{orb}$, $\overline{z}_{orb}$ $\left(\overline{y}_{orb}\right)$ is the mean value of the $z$ $\left(y\right)$ coordinate in the CM rest frame of the CM’s past orbit.
@@ Appendix B Estimation of radial velocities using machine learning
To complete the phase information for all the halo stars in our sample, we applied two machine learning techniques to assign radial velocities to those stars without such measurements.
@@ B.1 K-giant radial velocity
Our first sample consists of 245086 K-giant stars, of which 10989 have measured radial velocities. The last subsample was used to train a random forest regressor (RF) (Breiman, 2001). The chosen predictors for this study were the angular coordinates, proper motions, G magnitudes, BP and RP colours, distances to the Sun and the galactic centre, and the Galactocentric Cartesian coordinates. To prevent overfitting, a standard cross-validation analysis was performed. The RF hyper-parameters were tuned using GridSearchCV from the Scikit Learn library, resulting in the following values: [$n\_estimators=4900$, $random\_state=0$]
B.2 RR-Lyrae radial velocity
We observed a decrease in the number of measured radial velocities for RR-Lyrae stars, resulting in a drop in the fraction of measured radial velocities to (5510/67276), considering galactocentric distances between 10 to 100 kpc. To tackle this issue, we aimed to model the spatial distribution of radial velocities. To achieve this, we utilised normalising flows (NFs) (Durkan et al., 2019), as have been implemented by Crenshaw et al. (2021), to model the joint posterior probability of radial velocities and predictors using a subsample of the features described earlier. However, we limited the feature space to prevent any bias against less luminous stars. Therefore, we did not consider the magnitude and colours of stars as predictors in this case.
@@ B.3 Goodness of fit in radial velocity regression
R-Squared ($R^{2}$), or the coefficient of determination, is a statistical measure used to determine the proportion of variance in the dependent variable that can be explained by the independent variable in a regression model. The statistics were calculated for our two samples of stars, resulting in values of 0.86 for K-giant stars and 0.5 for RR-Lyrae ones, respectively. These values are comparable to those recently reported by Naik & Widmark (2023) using Bayesian neural networks, and provide us with a complete sample of halo stars in the phase space.
@@
Using this method, we can estimate the mean velocity field in the halo, which is necessary for the likelihood estimation of the subhalo mass of the LMC.
@@ Appendix C Validation of estimated distances for K-giant stars
To test the photometric distance obtained for the K giants, we computed the photometric distances of the globular clusters NGC7006, NGC5694, NGC2419, and NGC6229, and of the LMC, SMC, Carina, Draco, and Sculptor.
The observational data was extracted from Gaia Data Release 3 (Gaia Collaboration et al., 2023, 2016), using the option single object searcher to avoid field-contamination. Additionally, the data for the LMC and the SMC were selected on a circular disk in $(l,b)$ centred on the location of each MC, with a radius of $2^{\circ}$. For Carina, Draco, and Sculptor we selected data using an angular mask of three times the tidal radius reported in Drlica-Wagner et al. (2020a). We performed the data reduction or treatment indicated in the text for the K giant (that is, corrected for dust extinction, discarded all sources with $E(B-V)>0.3$, corrected the magnitudes, and performed the $3\sigma$ cut upon the corrected BP and RP flux excess factor ($C^{*}$)). Finally, we selected the giant branch and performed the cross-match with the K-giant catalogue given by Gaia (gaiadr3.astrophysical parameters). For the LMC and SMC, we also restricted the dataset in the reported Gaia parallax.
We calculated the photometric distance of each source and the mean value of each object. The results are shown in Fig. 5 and, as can be seen, the computed photometric distances and the values reported in the literature (Harris, 1996b; Drlica-Wagner et al., 2020b) are in good agreement.
Figure 5: Comparison between the photometric distances computed using Eq. (1) and the tabulated distances (Harris, 1996b; Drlica-Wagner et al., 2020b). Solid olive line: photometric distance equals tabulated distance. Dashed brown line: ±$\pm 10\%$ with respect to the one-to-one line. Solid purple line: ±$\pm 20\%$ with respect to the one-to-one line. The vertical lines are the statistical errors.